# Modeling Neural Activity with Conditionally Linear Dynamical Systems

### v1

@@ A.3 Correspondence between CLDS and Switching LDS
$\mathrm{cov}({\textnormal{a}}_{t},{\textnormal{a}}_{t+n})=\mathbb{E}\left[{\textnormal{a}}_{t}{\textnormal{a}}_{t+n}\right]-\mathbb{E}\left[{\textnormal{a}}_{t}\right]\mathbb{E}\left[{\textnormal{a}}_{t+n}\right]=\bm{a}^{\top}\left(P^{n}\Pi-\pi\pi^{\top}\right)\bm{a}$ (51)
@@
$\mathrm{cov}({\textnormal{a}}(t_{i}),{\textnormal{a}}(t_{j}))=k_{t}(t_{i},t_{j})=\bm{a}^{\top}\left(P^{\left|t_{j}-t_{i}\right|}\Pi-\pi\pi^{\top}\right)\bm{a}$ (52)
@@ Appendix B Experiments
B.1 Task-agnostic model implementations
We initialize observation matrices for all models (${\bm{C}}$ in (C)LDS models, log-rate decoder weights in LFADS) as the PCA principal axes in ${\mathbf{y}}$-space—that is, the top $D$ right singular vectors of the data—for each task.
@@ B.1 Task-agnostic model implementations
In all experiments, we assume that ${\mathbf{d}}({\bm{u}}_{t})=\bm{0}$, which forces the predicted firing rates, conditioned on ${\bm{u}}_{t}$, to lie in a $D$-dimensional space spanned by the columns of ${\mathbf{C}}({\bm{u}}_{t})$.
@@
We use the Jax implementation of LFADS available at https://github.com/google-research/computation-thru-dynamics/tree/master/lfads_tutorial. We choose the factor dimension to be the same as the latent dimension $D$ of the (C)LDS models and the inferred inputs to be of dimension $\left|\mathcal{U}\right|$, both following the (C)LDS models on any given experiment. The other components of the architecture are held fixed across all experiments:
@@ B.2 Synthetic experiment and parameter recovery
For the peaks $\xi_{i}$ spanning regularly the interval $[-\pi,\pi)$ and widths $\gamma$, the tuning curves in the rows of ${\mathbf{C}}$ for the synthetic experiment are defined as
@@
B.3 Pre-processing
@@
We considered neural activity from the “wake” period, binned in 50ms time-bins, then processed to firing rates and separated into 10s trials.
@@
We analyzed neural recordings of dorsal premotor cortex (PMd) in macaques performing center-out reaching task from N. Even-Chen, B. Sheffer et al. (2019). The experiments supported 3 different reach radii, and we selected only the middle reach radius at $8$cm. We were left with only the angular direction as the reach condition, over $N=16$ possible reach angles. We aligned all trials around the go-cue, selecting 200 ms before the go-cue and 300ms after. We binned the data into $5$ms bins, and performed Gaussian kernel smoothing with a standard deviation of 0.5 over bins.
B.4 Hyper-parameters
Throughout all experiments, we’ve set $L=5$ to balance expressivity and number of parameters. To select the other hyper-parameters of GP prior length-scale $\kappa$ and scale $\sigma$, we evaluated the various models on a held-out validation set. We plot in Fig. 6 our search over $\{\kappa,\sigma\}$ on the macaque center-out reaching data.

### v2

@@ A.3 Correspondence between CLDS and Switching LDS
$\mathbb{C}\mathrm{ov}({\textnormal{a}}_{t},{\textnormal{a}}_{t+n})=\mathbb{E}\left[{\textnormal{a}}_{t}{\textnormal{a}}_{t+n}\right]-\mathbb{E}\left[{\textnormal{a}}_{t}\right]\mathbb{E}\left[{\textnormal{a}}_{t+n}\right]=\bm{a}^{\top}\left(P^{n}\Pi-\pi\pi^{\top}\right)\bm{a},$
@@
$\mathbb{C}\mathrm{ov}({\textnormal{a}}(t_{i}),{\textnormal{a}}(t_{j}))=k_{t}(t_{i},t_{j})=\bm{a}^{\top}\left(P^{\left|t_{j}-t_{i}\right|}\Pi-\pi\pi^{\top}\right)\bm{a}.$ (28)
@@ Appendix B Experiments
B.1 Model implementations
We initialize observation matrices for all models ($\bm{C}$} in LDS, CLDS and gpSLDS models, log-rate decoder weights in LFADS) as the PCA principal axes in $\mathbf{y}$}-space—that is, the top $D$ right singular vectors of the data—for each task.
@@ B.1 Model implementations
In all experiments, we assume that ${\mathbf{d}}({\bm{u}}_{t})=\bm{0}$, which forces the predicted firing rates, conditioned on ${\bm{u}}_{t}$, to lie in a $N$-dimensional space spanned by the columns of ${\mathbf{C}}({\bm{u}}_{t})$.
gpSLDS
We implemented the gpSLDS [19] using its accompanying code at https://github.com/lindermanlab/gpslds. We used a latent dimension of $D=2$ in all experiments as we encountered significant OOM errors with higher dimensions. We used exactly the same inputs ${\bm{u}}_{1:T}$ as those provided to the (C)LDS models. We use $J=2$ and the feature transformation
$\bm{\phi}({\bm{x}})=\begin{bmatrix}1&{\bm{x}}_{1}&{\bm{x}}_{2}&{\bm{x}}_{1}^{2}&{\bm{x}}_{2}^{2}\end{bmatrix}^{\top}$ (30)
as to allow both linear (as in the rSLDS) and circular $\bm{\pi}({\bm{x}})$ partitions. We experimented with providing the same basis functions for the feature transformation as the ones provided to the CLDS but found that it did not provide any advantage. Other hyper-parameters we set to default provided values.
@@
We use the Jax implementation of LFADS available at https://github.com/google-research/computation-thru-dynamics/tree/master/lfads_tutorial, which is under a Apache-2.0 license. We choose the factor dimension to be the same as the latent dimension $D$ of the (C)LDS models and the inferred inputs to be of dimension $\left|\mathcal{U}\right|$, both following the (C)LDS models on any given experiment. The other components of the architecture are held fixed across all experiments:
@@ B.2 Synthetic experiment and parameter recovery
For the peaks $\xi_{i}$ spanning regularly the interval $\pi,\pi)$ and widths $\gamma$, the tuning curves in the rows of $\mathbf{C}$} for the synthetic experiment are defined as
@@
In Table 1, we report the results of our inference method on the first synthetic ring-attractor experiment (§3.2) for varying emission noise scale $R=\sigma^{2}_{R}I$, keeping all other experimental setup parameters exactly the same as in the original experiment and fixed. We see that even as the signal-to-noise ratio degrades and we reach a co-smoothing value of $0.21$, which is lower than our fit on the monkey data of §3.5, we still have decent $\mathbf{A}$} recovery and near-perfect $R$ recovery.
True noise log scale $\log\sigma_{R}$ $\bm{-2}$ $\bm{-1}$ $\bm{0}$ $\bm{1}$
Recovered $R$ log scale 1.97-1.97 0.98-0.98 $0.02$ $1.02$
${\mathbf{A}}(\cdot)$ recovery error $0.01$ $0.02$ $0.11$ $0.32$
Co-smoothing $R^{2}$ $0.99$ $0.94$ $0.68$ $0.21$
Table 1: Impact of observational noise on quality of fit. The “recovered $R$ scale” is the square-root of the matrix 22-norm of $R$. The ${\mathbf{A}}(\cdot)$ recovery error is the average in 2-norm error (between recovered and true) in eigenvalues over a regular grid of 50 angular conditions $\theta$. The co-smoothing $R^{2}$ is the average validation set $R^{2}$ on top-5-variance neurons, like in the main text.
B.3 Neural circuit ring attractor experiment
Let $y(\phi,t)$ be the firing-rate “bump” on angles $\phi\in[0,2\pi)$ with periodic boundary conditions, and let $v(t)=\dot{\theta}(t)$ be the angular velocity derived from the head-direction input. We consider its dynamics governed by the stochastic PDE
$dy(\phi,t)=\tau\left[-y(\phi,t)+f\big((w*y)(\phi,t)\big)+\kappa_{v}v(t)\,\partial_{\phi}y(\phi,t)\right]dt+\sigma dB(\phi,t)$
with
  • $f(\cdot)$ pointwise ReLU nonlinearity.
  • The recurrent drive is a circular convolution
    $(w*y)(\phi,t)\;=\;\int_{0}^{2\pi}w(\phi-\psi)\,y(\psi,t)\,d\psi,$
    with a rotationally symmetric kernel (in radians)
    $W(\Delta)=J_{e}\exp\left(-\tfrac{1}{2}(\Delta/\sigma_{e})^{2}\right)-J_{i}\exp\left(-\tfrac{1}{2}(\Delta/\sigma_{i})^{2}\right)$
    for $\Delta\in(-\pi,\pi]$ (wrapped), $\{J_{e},\sigma_{e}\},\{J_{i},\sigma_{i}\}$ parameters of excitation and inhibition.
  • The term $\kappa_{v}\,v(t)\,\partial_{\phi}y$ shifts the bump at speed proportional to $v(t)$, of parameter $\kappa_{v}$.
  • $B$ cylindrical Wiener process on the ring with covariance $\mathbb{E}[B(\phi,t)\,B(\phi^{\prime},t^{\prime})]=\delta(\phi-\phi^{\prime})\,\min(t,t^{\prime})$ (“white-noise” $dB/dt$).
  • $\tau>0$ a the time constant.
Numerical implementation
We discretize time to $\Delta t$ and head-direction to $\phi_{i}$, $i\in\{1,\dots,N\}$, $N=32$, regularly spanning $0,2\pi)$. Let $y(\phi_{i},n\Delta t)=[{\mathbf{y}}_{n}]_{i}$. Consider $W$ a circulant matrix implementing $w$, and $D$ a centered difference discretization of the derivative. Our update is
$\dot{\mathbf{y}}_{t}$ $=-{\mathbf{y}}_{t}+f(W{\mathbf{y}}_{t})+\kappa_{v}\,\theta_{t}\,D{\mathbf{y}}_{t}$
${\mathbf{y}}_{t+1}$ $={\mathbf{y}}_{t}+\tau\,\Delta t\,\dot{\mathbf{y}}_{t}+\sigma\,\sqrt{\Delta t}\,\eta,\quad\eta\sim\mathcal{N}(0,1)$
as the Euler–Maruyama discretization of the SPDE.
After a search over kernel hyperparameters, we found the best-performing model (Fig. 7) to encode a ring of fixed points that closely resembles the synthetic HD fixed points (§3.2), and with eigenvalues closer to the ones observed from the CLDS fits on the mouse ADn recordings (§3.4). The results make us confident that the CLDS can capture this nonlinear structure even under model mismatch.
Figure 7: Inferred CLDS fixed points (a), local dynamics (b) and eigenvalues of $\mathbf{A}$} (c) as a function of head-direction $\theta$ for the neural circuit ring attractor experiment. Hyperparameter values are $\{\sigma,\kappa\}=\{1.0,0.4\}$, obtained held-out $L$ over a grid search.
B.4 Data and Pre-processing
@@
We analyze neural recordings of antero-dorsal thalamic nucleus (ADn) from Ref. [32] of the mouse HD system in mice foraging in an open environment (Fig. 3a). The data was accessed through the pynapple (PYthon Neural Analysis Package, https://pynapple.org/, MIT License) package, with the data itself stored in NWB format on OSF at https://osf.io/jb2gd.
We considered neural activity from the “wake” period, binned in 50ms time-bins, then processed to firing rates by smoothing over a window of 4 bins, and separated into 10s trials.
@@
We analyzed neural recordings of dorsal premotor cortex (PMd) in macaques performing center-out reaching task from Ref. [33]—the associated article is under a CC BY 4.0 license.
The experiments supported 3 different reach radii, and we selected only the middle reach radius at 88cm. We were left with only the angular direction as the reach condition, over $N=16$ possible reach angles. We aligned all trials around the go-cue, selecting 200 ms before the go-cue and 300ms after. We binned the data into 55ms bins, and performed Gaussian kernel smoothing with a standard deviation of 0.5 over bins.
B.5 Hyper-parameters
Throughout all experiments, we’ve set $L=5$ to balance expressivity and number of parameters. While it is possible to treat $L$ as a hyperparameter tuned to improve performance, we propose to instead set $L$ to a large enough value that results in negligible error in the GP kernel approximation. As $L\to\infty$, we pay a larger computational price in fitting the model, but the statistical properties of the model are essentially unchanged because high-frequency basis functions have nearly zero amplitude. This idea is well established in GP regression literature [56].
To select the other hyper-parameters of GP prior length-scale $\kappa$ and scale $\sigma$, we evaluated the various models on a held-out validation set. We plot in Fig. 6 our search over $\{\kappa,\sigma\}$ on the macaque center-out reaching data.
Finally, a dimensionality of $D=2$ for the head-direction experiments is chosen throughout for interpretability, as the latent space is thought to encode head-direction. For the macaque center-out reaching data, we computed co-smoothing over $D\in\{3,5,10,15\}$ (Table 2) and selected $D=5$.
$\bm{D}$ $\bm{3}$ $\bm{5}$ $\bm{10}$ $\bm{15}$
CLDS $0.229$ $0.232$ $0.207$ $0.168$
LDS $0.211$ $0.193$ $0.196$ $0.170$
Table 2: Co-smoothing $R^{2}$ reconstruction of single held-out neurons from the test-set, over varying latent dimensionality $D$, averaged over two random seeds.
B.6 Compute Resources
We list below the compute resources used per experiment:
  1. 1.
    Synthetic HD experiment: Results were computed on a personal machine (Apple MacBook Pro, M2 Pro chip, 32G RAM). Inference converges in a few seconds.
  2. 2.
    Mouse HD experiment: Results were computed on a personal machine (Apple MacBook Pro, M2 Pro chip, 32G RAM). Inference converges in a few minutes.
  3. 3.
    Monkey-reaching experiment: Results were computed on an external cluster equipped with NVIDIA A100 GPUs. Fitting a single model typically used 60GB of GPU memory. All CLDS models were run strictly on CPU-only nodes (Intel Xeon Silver 4309Y Processor, 2.80 GHz, 8 cores/16 threads), using approximately 2 GB of RAM per run. Wall-clock times were between ten minutes and one hour.